OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -3/4 and b = -2.

OpenStudy (the_fizicx99):

y = -3/4x - 2 ? Is that what you meant?

OpenStudy (anonymous):

y = -3/4x - 2 --- add 3/4x to both sides 3/4x + y = -2 -- multiply by the LCD, which is 4 3x + 4y = -8 --- standard form

OpenStudy (anonymous):

do you have any questions ?

OpenStudy (anonymous):

no, thanks.